# DRExplainer: Quantifiable Interpretability in Drug Response Prediction with Directed Graph Convolutional Network

### v1

@@ 1 Introduction
The success of drug development significantly influences advancements in modern medicine [1]. However, substantial inter-individual variability in drug responses has led to uncertainties in therapeutic outcomes across clinical settings. In modern medical practice, accurately predicting an individual’s specific response while minimizing adverse reactions has become one of the key challenges in achieving personalized medicine and enhancing therapeutic efficacy [2]. Nevertheless, the field of drug response prediction encounters multifaceted challenges. The interplay of genetic diversity and the variation of environmental influences bring uncertainty to drug response prediction [3]. The metamorphosis of drug metabolism and the spectrum of mechanisms of action across individuals generate diverse effects from uniform drug doses, unraveling a series of responses. Coalesced with these genetic traits are the dynamic layers of gene expressions, whose complexity shapes the variability of drug responses.
@@
The field of graph-based learning has witnessed rapid-developed interest among researchers, leading to its rapid-developed application in the realm of bioinformatics [13, 14, 50]. Notably, these techniques have demonstrated remarkable utility in understanding complex biological systems. While graph-based learning methodologies have burgeoned across various domains [15, 16], the majority of approaches in drug response prediction have primarily operated within the confines of undirected bipartite networks. For example, Liu et al. utilized a hybrid graph convolutional network and multiple subnetworks for accurate prediction of cancer drug response [17], Liu et al. constructed a graph neural network with contrastive learning to enhance the generalization ability of drug response prediction [18], Zhu et al. proposed a framework consists of twin graph neural network and a similarity augmentation for drug response prediction [19]. Typically, a thorough investigation of drug response demands a separate analysis of sensitivity and resistance. Thus, the necessity to advance algorithms for directed bipartite networks becomes apparent.
@@
Furthermore, model interpretability has emerged as a critical concern. In clinical practice, stakeholders, including physicians and patients, require not only accurate predictions but also a clear understanding of how and why a model arrives at a specific prediction. Many high-performing deep learning models, such as deep neural networks, are often regarded as “black boxes”, impeding their interpretability [20]. Recently, several approaches have been proposed to explain the predictions based on graphs since graph neural networks have become increasingly popular [21, 22]. This interpretability improves their adoption and trustworthiness in critical medical decision-making. Current efforts to enhance the interpretability of drug response prediction models predominantly rely on algorithms designed for homogeneous networks [23]. For example, researchers explored methodologies for identifying and ranking the most influential features or nodes within homogeneous networks. Huang et al. designed a nonlinear model to explain the node in the subgraph, computing the k most representative features [24]. Vu et al. proposed a probabilistic graphical model to identify the crucial nodes and the dependencies of explained features in the form of conditional probabilities [25]. There are also explanation methods that identify exiting links in the network to explain the predicted links based on gradient or mutual information, such as GNNExplainer [26] and ExplaiNE [27]. Some studies delved into the identification of significant pathways and modules within homogeneous networks. However, there is a significant absence of interpretability methods tailored for directed bipartite networks. Additionally, the absence of standardized approaches that can evaluate the level of interpretability across various methodologies hampers the rigorous assessment of interpretability.
@@
In this study, we have delved into a novel drug response prediction approach that addresses the pressing challenges in the field of current drug response prediction. Our main contributions are summarized as follows:
@@ 2.1 Materials
Our study leveraged a combination of diverse data sources, carefully curated to ensure the comprehensiveness and accuracy of the information used in our research. This study mainly involves two databases, Genomics of Drug Sensitivity in Cancer (GDSC) [32] and Cancer Cell Line Encyclopedia (CCLE) [4]. CCLE stands as a rich repository of diverse cell lines’ multi-omics data, including genomics, transcriptomics, proteomics, and epigenomics. However, it’s essential to note that the CCLE dataset is relatively limited in terms of drug response data. In contrast, GDSC primarily focuses on capturing drug response information, offering a valuable collection of drug sensitivity and resistance profiles across a diverse spectrum of cancer cell lines. So, we harnessed the multi-omics profiles, which encompass gene expression, gene mutation, and copy number variation, from the CCLE dataset. Simultaneously, we collected response scores of IC50 values (natural log-transformed) from the GDSC dataset. In subsequent experiments, we used IC50 values as a measure of sensitivity.
@@
Following meticulous data collection, we obtained 986 cell lines, 358 drugs and a total of 295992 drug response data. During the sequencing and data collection of cell lines, it is common to encounter missing values. To optimize data richness, our focus was on including cell lines that possessed the three omics data types (i.e., gene expression, gene mutation and copy number variation). In order to thoroughly investigate the cancer-related mechanisms within drug response, we conducted preprocessing of the cell line omics data. Specifically, we retained the gene of cell line omics data mentioned in the COSMIC Cancer Gene Census [33]. And the three omics data types were ultimately represented as vectors with dimensions of 674, 689 and 480, respectively. Regarding drug data, we downloaded the SMILES strings of related drugs from PubChem [33] and DrugBank [34] datasets. It is crucial to acknowledge that not all drugs are encompassed within these two databases. A statistical summary of this dataset is presented in Table 1. After the collection of 157 drugs, we employed the ConvMolFeaturizer method in the DeepChem library [35] to compile each drug’s SMILES string into a molecular graph. In this representation, nodes and edges correspond to chemical atoms and bonds, respectively. This systematic conversion enables a structured and graph-based representation of each drug’s molecular composition.
@@
After rigorous preprocessing and data alignment, our experiment involved a curated dataset encompassing 477 cancer cell lines and 157 drugs. As mentioned above, we utilized IC50 values to measure the response between cell lines and drugs in this study. To categorize drug responses, we implemented two threshold values based on the IC50 values. Instances with IC50 values < -3 were classified as sensitive responses, whereas those with IC50 value > 3 were regarded as resistant. In a subsequent phase of our study, we compiled an expansive dataset featuring 17,071 drug responses, comprising 1,000 instances classified as sensitive and 16,071 as resistant. Specifically, we conducted the intricate drug response relationships within a directed graph framework, employing two adjacency matrices for representation. One adjacency matrix denoted sensitivity, where a value of 1 indicated a sensitive response, while the other matrix denoted resistance, with a value of 1 signifying a resistant response.
@@ 2.2.1 Node representation module
Inspired by the research of [17], we employed omics-specific neural network layers to integrate multi-omics data, aiming to achieve distinctive representation for each cancer cell line. A full depiction of a cancer cell line’s multi-omics data including gene expression, gene mutation and copy number variant feature vectors denoted by ${\boldsymbol{c}_{e}}$, ${\boldsymbol{c}_{m}}$, $\boldsymbol{c}_{n}$. These vectors were subsequently transformed into an F-dimensional vectorial embedding, as formulated in the equation:
@@ 2.2.2 Cell line and drug similarity
Building on the premise that cell lines with similar properties exhibit comparable drug sensitivities [37, 38], we regarded the similarity between cell lines and drugs as attributes of the cell lines and the drugs to infer drug sensitivity. Hence, we computed the similarity among cell lines and drugs, relying on the representations acquired in the section “Node representation module”. These similarities were quantified using cosine similarity, which performed pairwise global alignment among cell lines or drugs. The cell line-cell line similarity and the drug-drug similarity network were denoted by the matrices $\boldsymbol{S}^{c}=\left[S_{ij}^{c}\right]\in\mathbb{R}^{N_{c}\times N_{c}}$ and $\boldsymbol{S}^{d}=\left[S_{ij}^{d}\right]\in\mathbb{R}^{N_{d}\times N_{d}}$, respectively. Here, $S_{ij}^{c}$ and $S_{ij}^{d}$ quantified the pairwise similarities between cell lines $c_{i}$ and $c_{j}$, and between drugs $d_{i}$ and $d_{i}$. The total number of cell lines and drugs were denoted by $N_{r}$ and $N_{f}$. In this investigation, we employed cosine similarity to measure the similarity among cell lines or drugs, with the similarity score represented as $\cos s_{ij}$, which evaluates the degree of similarity between cell lines $c_{i}$ and $c_{i}$ or drugs $d_{i}$ and $d_{j}$. The cosine similarity was calculated as follows:
@@
where $\boldsymbol{x}_{\boldsymbol{i}}$ and $\boldsymbol{x}_{\boldsymbol{i}}$ represented the feature vectors of entity $i$ and $j$, respectively. To reduce the negative effect arising from apparent similarity data, we introduced a threshold, denoted by $\varphi$, to constrain the distribution of the similarity data. Specifically, we retained only similarity values where $\varphi>0.9$ for cell lines and $\varphi>0.88$ for drugs, thereby focusing on the most significant relationships. Hence, the similarity matrix can be obtained as follows:
@@ 2.2.4 DistMult decoder
Regarding the model’s overall architecture, DRExplainer can be characterized as $\Phi$ learning a conditional distribution $P_{\Phi}(r\mid\mathcal{A},\boldsymbol{X})$. As in previous work on factorization [40], we trained the model with negative samples. The positive samples in the training set were often randomly corrupted to generate negative samples. This approach will result in the generated negative sample being consistent or paradoxical with the positive sample. Inspired by Rossi et al. [41], we generated negative triples by treating the unutilized data in the data space as negative samples. To enhance model performance, we employed cross-entropy loss optimization, aiming to encourage the model to assign higher scores to observable triples compared to negative ones. The loss function was mathematically expressed as:
@@ 2.4 Constructing ground truths and evaluating
Figure 3: The AUC scores of DRExplainer in inductive capability learning.
@@ 3.1 Experimental design
Figure 3 delineated a comparative assessment of our model’s inductive capability across four distinct experimental tasks, which were designed to rigorously test the model’s predictive accuracy in various scenarios. Each task’s performance was illustrated via the Receiver Operating Characteristic (ROC) curve, providing insight into the model’s performance. TaskA adopted a traditional experimental setting, utilizing a test dataset composed of known drugs and cell lines to predict novel responses. And TaskA achieved the highest Area Under the Curve (AUC), indicating robust predictive ability within known entities. This setup served as a benchmark for the model’s performance under familiar conditions. In the analysis of inductive capability concerning cell lines and drugs, depicted in Task B and Task C, the performance of DRExplainer for cell lines exhibited a modest decline compared to drugs. This discrepancy may stem from the truth that cell lines share similar genetic information while chemical structures of drugs can be diversified. TaskD, while presenting the most challenging conditions, still showcased considerable predictive performance, albeit lower than the other tasks. Together, the AUC results across all tasks underscored the model’s robust inductive capability, highlighting its potential utility in real-world clinical settings.
@@ 3.2 Predicting method comparison
To verify the effectiveness of our model DRExplainer, we conducted comparative analyses against several established machine learning algorithms, including support vector machine (SVM), random forest (RF), and XGBoost. In order to illustrate the superiority of our model proposed in this study, we also benchmarked against three state-of-the-art (SOTA) algorithms: MOFGCN [11], DrugCell [31], and TGSA [19], and the latter two of which were extensively discussed in the literature [14]. Each algorithm underwent a five-fold cross-validation experiment on our dataset for a comprehensive performance assessment.
Figure 4 presented the comparative results of machine learning approaches, focusing on the Area Under the Precision-Recall Curve (AUPR) and the Area Under the Receiver Operating Characteristic Curve (AUC). To elucidate the findings in the figure, we reported the results of the best performing fold. The outcomes demonstrated that our model consistently surpassed the comparative algorithms, achieving the highest scores in both AUPR and AUC metrics. The performance analysis revealed that conventional machine learning algorithms frequently encountered difficulties when processing a large amount of data and complex associations [45], which contributed to their suboptimal performance relative to our model. To rigorously assess the efficacy of our model, we compared our method with the following SOTA methods.
@@
Figure 5 delineated the performance across different metrics, providing a comprehensive evaluation of our model’s capabilities. The AUC served as a global measure of model performance, independent of any specific classification threshold, meanwhile, the AUPR provided a more detailed assessment, particularly beneficial for imbalanced datasets, by focusing on the precision-recall balance across varying thresholds. Both metrics were essential for a comprehensive evaluation of classification models, with higher values signifying superior class discrimination and predictive accuracy. In terms of both AUC and AUPR, our model exhibited outstanding performance, outperforming other SOTA models. Specifically, it exceeded the AUC scores of MOFGCN, TGSA, and DrugCell by 13.76%, 8.36%, and 4.34%, respectively, and surpassed them in AUPR scores by 13.76%, 8.36%, and 4.34%, respectively. These results highlighted the exceptional performance and robustness of our model in classification tasks.
In addition, our model showcased remarkable specificity, an essential trait in analyses where negative samples predominated over positive ones. Nonetheless, it was important to acknowledge that, despite its overall superiority, our model did not attain the highest F1 score when compared to MOFGCN and DrugCell. This divergence was ascribed to MOFGCN and DrugCell’s higher recall, which cumulatively contributed to their augmented F1 score. And our model’s ability to tackle complex biological problems, as mentioned above, was noteworthy. DRExplainer extended beyond the common prediction of sensitivity to accurately identify resistance and non-significant relationships, demonstrating its contribution to the field of drug response research.

### v2

@@ 1 Introduction
The success of drug development significantly influences advancements in modern medicine [1]. In modern medical practice, accurately predicting an individual’s specific response while minimizing adverse reactions has become one of the key challenges in achieving personalized medicine and enhancing therapeutic efficacy [2]. Predicting the response of cancer cell lines to therapeutic drugs is a cornerstone of precision medicine, where treatments are tailored to individual patients based on their unique genetic profiles. This is particularly crucial in oncology, where inter-individual variability in drug responses can lead to ineffective treatments or severe adverse effects. Nevertheless, the field of drug response prediction encounters multifaceted challenges. The interplay of genetic diversity and the variation of environmental influences bring uncertainty to drug response prediction [3]. Coalesced with these genetic traits are the dynamic layers of gene expressions, whose complexity shapes the variability of drug responses.
@@
The field of graph-based learning has witnessed rapid-developed interest among researchers, leading to its rapid-developed application in the realm of bioinformatics [14, 15, 16, 17]. Notably, these techniques have demonstrated remarkable utility in understanding complex biological systems. While graph-based learning methodologies have burgeoned across various domains [18, 19, 20], the majority of approaches in drug response prediction have primarily operated within the confines of undirected bipartite networks. For example, Liu et al. utilized a hybrid graph convolutional network and multiple subnetworks for accurate prediction of cancer drug response [21], Liu et al. constructed a graph neural network with contrastive learning to enhance the generalization ability of drug response prediction [22], Zhu et al. proposed a framework consists of twin graph neural network and a similarity augmentation for drug response prediction [23]. Typically, a thorough investigation of drug response demands a separate analysis of sensitivity and resistance. Thus, the necessity to advance algorithms for directed bipartite networks becomes apparent.
@@
Furthermore, model interpretability has emerged as a critical concern. Despite significant advances, existing predictive models often lack the interpretability required for clinical adoption, limiting their utility for medical decision-making. In clinical practice, stakeholders, including physicians and patients, require not only accurate predictions but also a clear understanding of how and why a model arrives at a specific prediction. Many high-performing deep learning models, such as deep neural networks, are often regarded as “black boxes”, impeding their interpretability [24]. Recently, several approaches have been proposed to explain the predictions based on graphs since graph neural networks have become increasingly popular [25, 26]. This interpretability improves their adoption and trustworthiness in critical medical decision-making. Current efforts to enhance the interpretability of drug response prediction models predominantly rely on algorithms designed for homogeneous networks [27]. For example, researchers explored methodologies for identifying and ranking the most influential features or nodes within homogeneous networks. Huang et al. designed a nonlinear model to explain the node in the subgraph, computing the k most representative features [28]. Vu et al. proposed a probabilistic graphical model to identify the crucial nodes and the dependencies of explained features in the form of conditional probabilities [29]. There are also explanation methods that identify exiting links in the network to explain the predicted links based on gradient or mutual information, such as GNNExplainer [30] and ExplaiNE [31]. Some studies delved into the identification of significant pathways and modules within homogeneous networks. However, there is a significant absence of interpretability methods tailored for directed bipartite networks. Additionally, the absence of standardized approaches that can evaluate the level of interpretability across various methodologies hampers the rigorous assessment of interpretability.
@@
Existing drug response prediction methods face critical challenges, including limited interpretability, dependence on undirected networks, and the lack of standardized evaluation approaches for interpretability. In this study, we presented DRExplainer, a novel approach that bridges these gaps by integrating a directed graph convolutional framework with a dedicated interpretability algorithm. Furthermore, we constructed a ground truth benchmark dataset to quantitatively evaluate the interpretability of the model, which sets it apart from existing approaches. Our main contributions are summarized as follows:
@@ 2.1 Materials
Our study leveraged a combination of diverse data sources, carefully curated to ensure the completeness and precision of the information used in our research. This study mainly involves two databases, Genomics of Drug Sensitivity in Cancer (GDSC) [36] and Cancer Cell Line Encyclopedia (CCLE) [4]. CCLE stands as a rich repository of multi-omics data from diverse cell lines, including genomics, transcriptomics, proteomics, and epigenomics. However, it’s essential to note that the CCLE dataset is relatively limited in terms of drug response data. In contrast, GDSC primarily focuses on capturing drug response information, offering a valuable collection of drug sensitivity and resistance profiles across a diverse spectrum of cancer cell lines. So, we harnessed the multi-omics profiles, which encompass gene expression, gene mutation, and copy number variation, from the CCLE dataset. Simultaneously, we collected response scores of IC50 values (natural log-transformed) from the GDSC dataset. In subsequent experiments, we used IC50 values as a measure of sensitivity.
@@
Following meticulous data collection, we obtained 986 cell lines, 358 drugs and a total of 295992 drug response data. During the sequencing and data collection of cell lines, it is common to encounter missing values. To ensure data quality and completeness, we focused on cell lines that possessed the three omics data types (i.e., gene expression, gene mutation and copy number variation). In order to thoroughly investigate the cancer-related mechanisms within drug response, we preprocessed the cell line omics data. Specifically, we retained the gene of cell line omics data listed in the COSMIC Cancer Gene Census [37]. This preprocessing resulted in feature vectors of dimensions 674, 689, and 480 for gene expression, gene mutation, and copy number variation, respectively. For drug data, we obtained the SMILES strings of related drugs from PubChem [37] and DrugBank [38] datasets. It is crucial to acknowledge that not all drugs are encompassed within these two databases. Table 1 presents a statistical overview of the collected dataset. After the collection of 157 drugs, we employed the ConvMolFeaturizer method in the DeepChem library [39] to compile each drug’s SMILES string into a molecular graph. In this representation, nodes and edges correspond to chemical atoms and bonds, respectively. This systematic conversion enables a structured and graph-based representation of each drug’s molecular composition.
@@
After rigorous preprocessing and data alignment, the final curated dataset encompasses 477 cancer cell lines and 157 drugs. As mentioned above, we utilized IC50 values to measure the response between cell lines and drugs in this study. To categorize drug responses, we implemented two threshold values based on IC50 values. Instances with IC50 values < -3 were classified as sensitive responses, while those with IC50 value > 3 were regarded as resistant. In a subsequent phase of our study, we compiled an expansive dataset featuring 17,071 drug responses, comprising 1,000 instances classified as sensitive and 16,071 as resistant. Specifically, we conducted the intricate drug response relationships within a directed graph framework, employing two adjacency matrices for representation. One adjacency matrix denoted sensitivity, where a value of 1 indicated a sensitive response, while the other matrix denoted resistance, with a value of 1 signifying a resistant response.
@@ 2.2.1 Node representation module
Inspired by the research of [21], we implemented omics-specific neural network layers to integrate multi-omics data, aiming to achieve distinctive representation for each cancer cell line. A full depiction of a cancer cell line’s multi-omics data including gene expression, gene mutation and copy number variant feature vectors denoted by ${\boldsymbol{c}_{e}}$, ${\boldsymbol{c}_{m}}$, $\boldsymbol{c}_{n}$. These vectors were subsequently transformed into an F-dimensional vectorial embedding through the following equation:
@@ 2.2.2 Cell line and drug similarity
Building on the premise that cell lines with similar properties exhibit comparable drug sensitivities [41, 42], we incorporated cell line and drug similarity as key attributes for drug sensitivity inference. Hence, we computed the similarity among cell lines and drugs, relying on the representations obtained in the section “Node representation module”. These similarities were quantified using cosine similarity, which performed pairwise global alignment among cell lines or drugs. The cell line-cell line similarity and the drug-drug similarity network were denoted by the matrices $\boldsymbol{S}^{c}=\left[S_{ij}^{c}\right]\in\mathbb{R}^{N_{c}\times N_{c}}$ and $\boldsymbol{S}^{d}=\left[S_{ij}^{d}\right]\in\mathbb{R}^{N_{d}\times N_{d}}$, respectively. Here, $S_{ij}^{c}$ and $S_{ij}^{d}$ quantified the pairwise similarities between cell lines $c_{i}$ and $c_{j}$, and between drugs $d_{i}$ and $d_{i}$. The total number of cell lines and drugs were denoted by $N_{r}$ and $N_{f}$. In this investigation, we employed cosine similarity to measure the similarity among cell lines or drugs, with the similarity score represented as $\cos s_{ij}$, which evaluates the degree of similarity between cell lines $c_{i}$ and $c_{i}$ or drugs $d_{i}$ and $d_{j}$. The cosine similarity was calculated as follows:
@@
where $\boldsymbol{x}_{\boldsymbol{i}}$ and $\boldsymbol{x}_{\boldsymbol{i}}$ represented the feature vectors of entity $i$ and $j$, respectively. To mitigate noise from weak similarities, we introduced a threshold, denoted by $\varphi$, to constrain the distribution of the similarity data. Specifically, we retained only similarity values where $\varphi>0.9$ for cell lines and $\varphi>0.88$ for drugs, thereby focusing on the most significant relationships. Hence, the similarity matrix can be obtained as follows:
@@ 2.2.4 DistMult decoder
Regarding the model’s overall architecture, DRExplainer can be characterized as $\Phi$ learning a conditional distribution $P_{\Phi}(r\mid\mathcal{A},\boldsymbol{X})$. As in previous work on factorization [44], we trained the model with negative samples. The positive samples in the training set were often randomly corrupted to generate negative samples. This approach will result in the generated negative sample being consistent or paradoxical with the positive sample. Inspired by these investigations [45, 46, 47], we generated negative triples by treating the unutilized data in the data space as negative samples. To enhance model performance, we employed cross-entropy loss optimization, aiming to encourage the model to assign higher scores to observable triples compared to negative ones. The loss function was mathematically expressed as:
@@ 2.4 Constructing ground truths and evaluating
Figure 3: The AUC and AUPR scores of DRExplainer in inductive capability learning.
@@ 3.1 Experimental design
Figure 3 delineated a comparative assessment of our model’s inductive capability across four distinct experimental tasks, which were designed to rigorously test the model’s predictive accuracy in various scenarios. Each task’s performance was illustrated via the Receiver Operating Characteristic (ROC) curve Precision-Recall Curve, providing insight into the model’s performance. TaskA adopted a traditional experimental setting, utilizing a test dataset composed of known drugs and cell lines to predict novel responses. And TaskA achieved the highest Area Under the Curve (AUC), indicating robust predictive ability within known entities. This setup served as a benchmark for the model’s performance under familiar conditions. In the analysis of inductive capability concerning cell lines and drugs, depicted in Task B and Task C, the performance of DRExplainer for cell lines exhibited a modest decline compared to drugs. This discrepancy may stem from the truth that cell lines share similar genetic information while chemical structures of drugs can be diversified. TaskD, while presenting the most challenging conditions, still showcased considerable predictive performance, albeit lower than the other tasks. Together, the results across all tasks underscored the model’s robust inductive capability, highlighting its potential utility in real-world clinical settings.
@@ 3.2 Predicting method comparison
To verify the effectiveness of our model DRExplainer, we conducted comparative analyses against several established machine learning algorithms, including support vector machine (SVM), random forest (RF), and XGBoost. In order to illustrate the superiority of our model proposed in this study, we also benchmarked against three state-of-the-art (SOTA) algorithms: MOFGCN [11], NIHGCN [51], DrugCell [35], and TGSA [23], and the latter two of which were extensively discussed in the literature [15]. Notably, MOFGCN, TGSA and NIHGCN are graph-based learning approaches widely applied in bioinformatics. Each algorithm underwent a five-fold cross-validation experiment on our dataset for a comprehensive performance assessment.
Figure 4 presented the comparative performance of various machine learning approaches, evaluated using the Area Under the Precision-Recall Curve (AUPR) and the Area Under the Receiver Operating Characteristic Curve (AUC). To ensure a fair and consistent comparison, these experiments were conducted using the feature derived from the Section ”Node representation module”. These features include multi-omics and molecular-level data, which are widely utilized in drug response prediction tasks. To elucidate the findings in the figure, we reported the results of the best performing fold. The results demonstrated that our model consistently surpassed the baselines algorithms, achieving the highest scores for both AUPR and AUC metrics. Further performance analysis revealed that conventional machine learning algorithms frequently encountered difficulties when processing a large amount of data and complex associations, as previously noted by Wang et al. [52], which contributed to their suboptimal performance relative to our model. More detailed experiments about machine learning algorithms are provided in the Supplementary material. To rigorously assess the efficacy of our model, we compared our method with the following SOTA methods.
@@
  • 4.
    NIHGCN leveraged an innovative heterogeneous graph convolutional network architecture, combining parallel graph convolution and neighborhood interaction layers to achieve precise predictions of anticancer drug responses.
Figure 5 delineated the performance across different metrics, providing a comprehensive evaluation of our model’s capabilities. The AUC served as a global measure of model performance, independent of any specific classification threshold, meanwhile, the AUPR provided a more detailed assessment, particularly beneficial for imbalanced datasets, by focusing on the precision-recall balance across varying thresholds. Together, these metrics were essential for a comprehensive evaluation of classification models, with higher values signifying superior class discrimination and predictive accuracy. In terms of both AUC and AUPR, our model exhibited outstanding performance, outperforming other SOTA models. Specifically, it exceeded the AUC scores of MOFGCN, TGSA, DrugCell, and NIHGCN by 13.76%, 8.36%, 4.34%, and 26.67%, respectively, and achieved similar margins of improvement in AUPR scores. These results highlighted the effectiveness and robustness of our model in drug response tasks.
In addition, our model showcased remarkable specificity, an essential trait in analyses where negative samples predominated over positive ones. Nonetheless, it was important to acknowledge that, despite its overall superiority, our model did not attain the highest F1 score when compared to MOFGCN and DrugCell. This divergence was ascribed to MOFGCN and DrugCell’s higher recall, which cumulatively contributed to their augmented F1 score. And our model’s ability to tackle complex biological problems, as mentioned above, was noteworthy. Moreover, we also conducted the inductive learning experiments of the SOTA methods and the results were provided in the Supplementary material. DRExplainer extended beyond the common prediction of sensitivity to accurately identify resistance and non-significant relationships, demonstrating its contribution to the field of drug response research.